OpenStudy (jpsmarinho):

How to inverse the function: Arcsin(1/2) = pi/6 ?? If possible, pass to pass ^^

OpenStudy (anonymous):

what's Arcsen?

OpenStudy (jpsmarinho):

Sorry, I'm Brazilian. Is arcsin (sin^(-1))

OpenStudy (anonymous):

yes.. Arcsin(1/2) = pi/6 is the same as saying: sin(pi/6) = 1/2 ....:)

OpenStudy (jpsmarinho):

But I know this ^^ . My question is how to conclude this pass by pass?

OpenStudy (jpsmarinho):

Indeed, what I want to know is how to proof that?

OpenStudy (jpsmarinho):

(Correction)Prove it in the place of proof that